-2(3s-4)-8=-2(7s+4)-3
Please help me solve this

Answer by HyperBrain(694)

-2(3s-4)-8=-2(7s+4)-3
(-6s+8)-8=(-14s-8)-3
-6s=-14s-11
8s=-11
s=-11/8
Power up,
HyperBrain!
Stop this mess! Click here